OpenStudy (anonymous):

Solve the following system of equations using the substitution method x+3y=11 xy=8

OpenStudy (anonymous):

x = 11-3y (11-3y)y=8 11y-3y^2-8=0 3y^2-11y+8=0 y = 2 and 2 thirds or y = 1 no plug it into the x = 11 -3y x = 3 x = 8 (3,2 and 2 thirds) (8,1)